OpenStudy (anonymous):

Sarah went to an electronics store, purchased 3 CDs and 5 DVDs, and spent $137. Michelle went to the same store, bought 2 CDs and 7 DVDs, and spent $161. How much does each CD sell for at the store?

OpenStudy (insa):

Let the cds= x then the dvds= y 3x+5y=137______ 1 2x+7y= 161__________2 2(3x+5y)= 2(137) 6x+10y= 274________ 3 3(2x+7y)=3(161) 6x+21y= 483_________4 Subtracting equation 3 and 4 6x+10y=274 6x+21y=483 - - - 11y = 209

OpenStudy (insa):

y= 209/11 y=19 3x+5(19)= 137 3x=137-95 3x=42 x=42/3 x=14 So, solution are= dvds=y=$19 cds=x=$14

OpenStudy (insa):

check= 2x+7y=161 2(14)+7(19)= 161 28+133=161 161=161